# Larimar: Large Language Models with Episodic Memory Control

### v1

@@ Appendix C ZsRE Single Fact Editing Results
Larimar demonstrates effective editing and paraphrasing on ZsRE, with comparable or slightly lower performance in the Neighborhood category, maintaining consistent results across GPT-2 and GPT-J decoders, underscoring its model-agnostic editing capabilities.
Editor Edit Success Paraphrase Neighborhood GPT-2 XL 22.2 21.3 24.2 FT 99.6 82.1 23.2 FT+L 92.3 47.2 23.4 KE 65.5 61.4 24.9 KE-zsRE 92.4 90.0 23.8 MEND 75.9 65.3 24.1 MEND-zsRE 99.4 99.3 24.1 ROME 99.8 88.1 24.2 Larimar-1.3B 98.1 81.6 19.7 GPT-J 26.4 25.8 27.0 ROME 99.8 95.9 27.2 Larimar-6B 94.5 70.4 25.1
@@ Appendix D Additional Counterfact Batch editing Results
Figure 4: Batch editing on CounterfFct dataset. Baseline performances are taken from [Meng et al., 2023]. Green: MEMIT, Orange: ROME, Magenta: MEND, Black: Larimar-6B.
@@ Appendix E Additional Experimental Details
For the sequential editing experiments reported in Table 3 and Figure 3, we set $K=1000$ and use a fixed reference memory $\mathbf{M}^{(\rm ref)}$ (see section 3) to compute reading and writing weights.
@@
For Figure 3, the reference memory is constructed by encoding the first prompt for each of the 1000 unique facts (among the several rephrasings in the edit set which are written to memory) and placing it in a single row in $\mathbf{M}^{(\rm ref)}$. Thus, when querying memory with an encoded rephrased prompt $\mathbf{z}$ in Eq. (7), if $\mathbf{z}$ is closest to the row $k$ in $\mathbf{M}^{(\rm ref)}$ corresponding to the same fact, the key vector element $w_{k}$ will be largest for this element, and suppressed for other memory locations. (We use $\alpha=10^{-3}$ to strongly suppress more distant encodings in the reference memory. Empirically, we found that that the nearest-neighbor encoding picked out by Eq. (7) with small $\alpha$ is usually the encoded prompt for the same fact, with lower F1 scores occurring mainly in cases where the nearest-neighbor row in $\mathbf{M}^{(\rm ref)}$ corresponds to a different fact.) We found that computing reading and writing weights as in [Pham et al., 2021], $\mathbf{w}=\mathbf{z}(\mathbf{M}^{(\rm ref)})^{\dagger}$, was not as effective with rephrased facts (Figure 3 and Table 13) unless the number of rephrasings per fact was relatively large.
@@
In Figure 5, we compare different variants of Larimar, on the same task as shown in Figure 3. Relative to the Gaussian convolution method of Eq. (7), computing reading and writing weights with the reference memory matrix pseudoinverse, $\mathbf{w}=\mathbf{z}(\mathbf{M}^{(\rm ref)})^{\dagger}$ performed well on a dataset of 511 ZsRE facts and $\approx 20$ phrasings per fact, but significantly worse on a dataset of $1000$ ZsRE with $10$ phrasings per fact. (We hypothesize that Eq. (7) is more effective at finding a nearby rephrase encoding for the same fact when there are only one or a few paraphrases available in the data.)
@@
Figure 5: Mean F1 score of Larimar, comparing different choices for computing reading and writing weights – the Gaussian convolution in Eq. (7) and the pseudoinverse method of [Pham et al., 2021] – on held-out sets of unseen rephrasings from ZsRE over a sequence of 3000 edits. (Black curves are shown in Figure 3 in the main text.)

### v2

@@ Appendix C ZsRE single fact editing experiments and ablations
The test-time augmentation of memory with additional paraphrases boosts generalization from 70.4% to 82.2%, when two rephrases written in memory. In absence of scope detector, the same approach of augmenting memory with two additional rephrases provide an additional 1-2% increase in generalization, whereas neighborhood specificity is affected significantly, irrespective of the dataset.
Editor Edit Success Paraphrase Neighborhood GPT-2 XL 22.2 21.3 24.2 FT 99.6 82.1 23.2 FT+L 92.3 47.2 23.4 KE 65.5 61.4 24.9 KE-zsRE 92.4 90.0 23.8 MEND 75.9 65.3 24.1 MEND-zsRE 99.4 99.3 24.1 ROME 99.8 88.1 24.2 Larimar-1.3B 98.1 81.6 19.7 GPT-J 26.4 25.8 27.0 ROME 99.8 95.9 27.2 Larimar-6B 94.5 70.4 25.1 Larimar-6B + 2 rephrases 94.5 82.2 25.1
@@
zSRE results Edit Success Paraphrase Neighb Larimar (W/ Scope) + 1 rephrase 91.8 77.9 25.1 Larimar (W/ Scope) + 2 rephrase 91.8 82.2 25.1 Larimar (W/O Scope) + 1 rephrase 92.1 79.4 8.6 Larimar (W/O Scope) + 2 rephrase 92.1 83.4 8.5
Table 13: Knowledge editing generalization results with and without additional rephrases and scope detectors on zSRE.
@@ Appendix D Additional Counterfact Batch editing Results
Figure 3: Batch editing on CounterFact dataset. Baseline performances are taken from [Meng et al., 2023]. Green: MEMIT, Orange: ROME, Magenta: MEND, Black: Larimar-6B.
@@
Figure 4: Batch editing on CounterFact dataset with different memory slot size $K$.
We use a memory of size $K\times C$, where K=512 and C=768 throughout the manuscript, unless otherwise stated. For the sequential editing experiments, where there were 1000 facts to be stored in memory, $K$ was set to 1000. Figure 4 shows the edit performance change as a function of the number of updates with different memory sizes. Results suggest that for $N=K=512$, rewrite accuracy is  99%. For $K=N>512$, the rewrite accuracy is slightly lower  94%, likely because Larimar was trained with K=512. For $N>K$, where K is smaller than 512, we find rewrite accuracy to be around 80% if $N=2K$ and around 54% if $N=4K$.
@@ Appendix E Additional Experimental Details
For the sequential editing experiments reported in Table 3 and Figure 2 (b), we set $K=1000$ and use a fixed reference memory $\mathbf{M}^{(\rm ref)}$ (see section 3) to compute reading and writing weights.
@@
For Figure 2 (b), the reference memory is constructed by encoding the first prompt for each of the 1000 unique facts (among the several rephrasings in the edit set which are written to memory) and placing it in a single row in $\mathbf{M}^{(\rm ref)}$. Thus, when querying memory with an encoded rephrased prompt $\mathbf{z}$ in Eq. (7), if $\mathbf{z}$ is closest to the row $k$ in $\mathbf{M}^{(\rm ref)}$ corresponding to the same fact, the key vector element $w_{k}$ will be largest for this element, and suppressed for other memory locations. (We use $\alpha=10^{-3}$ to strongly suppress more distant encodings in the reference memory. Empirically, we found that that the nearest-neighbor encoding picked out by Eq. (7) with small $\alpha$ is usually the encoded prompt for the same fact, with lower F1 scores occurring mainly in cases where the nearest-neighbor row in $\mathbf{M}^{(\rm ref)}$ corresponds to a different fact.) We found that computing reading and writing weights as in [Pham et al., 2021], $\mathbf{w}=\mathbf{z}(\mathbf{M}^{(\rm ref)})^{\dagger}$, was not as effective with rephrased facts (Figure 2 (b) and Table 14) unless the number of rephrasings per fact was relatively large.
@@
In Figure 5, we compare different variants of Larimar, on the same task as shown in Figure 2 (b). Relative to the Gaussian convolution method of Eq. (7), computing reading and writing weights with the reference memory matrix pseudoinverse, $\mathbf{w}=\mathbf{z}(\mathbf{M}^{(\rm ref)})^{\dagger}$ performed well on a dataset of 511 ZsRE facts and $\approx 20$ phrasings per fact, but significantly worse on a dataset of $1000$ ZsRE with $10$ phrasings per fact. (We hypothesize that Eq. (7) is more effective at finding a nearby rephrase encoding for the same fact when there are only one or a few paraphrases available in the data.)
@@
Figure 5: Mean F1 score of Larimar, comparing different choices for computing reading and writing weights – the Gaussian convolution in Eq. (7) and the pseudoinverse method of [Pham et al., 2021] – on held-out sets of unseen rephrasings from ZsRE over a sequence of 3000 edits. (Black curves are shown in Figure 2 (b) in the main text.)